The sum of two numbers is 55. The larger number is 7 more than twice the smaller number. what are the numbers?
Answer by wuwei96815(245)

x=one of the numbers
2x+7=the other number

x + (2x+7) = 55
3x + 7 = 55
3x = 55-7 = 48
x = 48/3 = 16
2x+7 = 2(16)+7 = 39

Did you follow the solution?